FIND THE DISTANCE BETWEEN THE POINTS A(4,-3) AND B(-4,3).
Answer by checkley75(3666)

the didtance between these two points (4,-3)& (-4,3)is
(4+4)^2+(3+3)^2=d^2
8^2+6^2=d^2
64+36=d^2
100=d^2\d=sqrt100
d=10 answer for the distance between these two points